dontsaymyname:

Solve the following quadratic equation by completing the square. 3x^2 -36x +96 =0

darkknight:

completing the square you have something in this form \[ax^2+bx+c = 0\] subtract c from both sides \[ax^2+bx = -c\] add (\[(b/2)^2\] to both sides so you have \[ax^2+bx+(b/2)^2=-c\] then you factor the left side of the eqn and solve from there

darkknight:

so for an example we take this 3x^2 -36x +96 =0 we want in form ax^2+bx+c so we divide 3x^2-36x+96 by 3 so we have \[x^2-12x+32=0\] (we just divided the whole eqn by 3) bring c to the other side \[x^2-12x=-32\] now b = -12 \[(b/2)^2=(-12/2)^2\] =36 so add 36 to both sides \[x^2-12x+36=-32+36\] \[x^2-12x+36=4\] factor left side of eqn into \[(x-6)^2=4\] take positive/negative square root of 4 \[x-6 =2 \] and \[x-6 =-2 \] x = 8 and x = 4 plug into the original function to see if it works

dontsaymyname:

TY :)

darkknight:

np